# skipabove=0.5skipbelow=0.5align=center

### v1

@@ IV-B Interaction Data – Security Signals and Metrics Definitions
Our system computes reputation from an actor’s historical interactions within the ecosystem. We categorize these interactions by security signals and quantify each signal using measurable metrics.
@@
    Security standards and recommendations: We consulted frameworks like the SLSA security framework [57], the CNCF software supply chain security guidelines [58], the NIST SSDF security framework [59], NIST SP 800-204D [60], Openssf S2C2F [61], and the CIS Software Supply Chain Security Guide[62]. Common recommendations across these sources were used to ensure the selection of well-established security practices.
  2. 2.
    Available security tools in the OSS ecosystem: We identified security tools available through GitHub’s user interface and API, which reflect the security capabilities easily available to contributors on the platform.
We grouped the resulting metrics into seven categories – Security Signals, based on contributors’ security tool usage and vulnerability management practices. A summary appears in  Table I. This resulted in seven security signals
IV-C Trust Engine – Reputation Computation
In this section, we outline the core functionality of the trust engine’s reputation score computation (see Figure 1). First, the interaction data formatter and security signal scorer extracts each contributor’s ecosystem interaction history and quantify each signal according to the metrics in Table II. These base measurements capture a user’s adherence to the defined security signals and their contribution patterns.
Next, ARMS evaluates each contributor’s reputation. To account for the risk posed by a contributor’s activities within the ecosystem, we refine the initial signal scores by incorporating:
Impact Score – (Weightage Signals Table I (W1-W3)):
  1. 1.
    Package Usage Weighting: Projects with no users (e.g., private repositories or those with zero forks/downloads) should not contribute to the user’s evaluation score, as they present minimal supply chain risk.
  2. 2.
    Community Tenure Weighting  [51]: This ensures that newer users do not unduly influence their reputation score, addressing concerns like those in the ”XZ utils” situation [63], where malicious activity was introduced by new members.
  3. 3.
    Centrality Score Weighting [18]: Centrality—an actor’s level of connectedness within the ecosystem based on contribution activities—can be crucial for scaling reputation. We suggest evaluating not only the number of connections (edges) but also the time taken to establish them.
Benchmark Score:
  1. 1.
    Combination of multiple security metrics/Composite Scoring: A holistic evaluation will aggregate a user’s performance across various security metrics, offering a comprehensive view of their adherence to security practices. This approach minimizes false positives and negatives by drawing from multiple metrics, as seen in the OpenSSF scorecard  [64]. Each metric’s weighting will be based on its effectiveness analysis section V-A.
  2. 2.
    Ecosystem Benchmarks: User performance will be benchmarked against the average or median reputation scores across the ecosystem to provide a relative measure of security adherence.
V Design of Experiments for ARMS
V-A Security Metrics Effectiveness Evaluation – Quasi-Experimental Analyses
To evaluate the utility of the proposed security signals, we propose the following quasi-experimental studies to determine whether the signals reliably predict poor security practices.
  1. 1.
    Inter-metric Relationship Study: Compare OSS projects that adopted a given security practice (e.g., branch protection) to matched control projects that did not, using a difference‐in‐differences design. We will analyze the results with regression models incorporating project and time fixed effects and interaction terms for the paired metrics to isolate their combined impact on security outcomes.
    • Data: Historical GitHub records for projects within the same domain, similar size, and activity levels.
    • Outcomes: Changes in vulnerability incidence (Signals 1–2) and shifts in contributor behaviors (Signals 3–7).
    • Ethics: Uses only publicly available data, avoiding additional data collection or privacy concerns.
  2. 2.
    Retrospective Incident Prediction Study: Identify projects that experienced known supply-chain incidents (e.g., npm/Event-Stream, ua-parser-js) and compare their pre-incident signal profiles against similar “clean” projects.
    • Data: Metric values for each project’s maintainers before the incident.
    • Analysis: Logistic regression (with interaction terms) to determine which signals best predict incident occurrence.
V-B User Behavioural Studies – Surveys & Interviews
  1. 1.
    Collaborator Vetting Study: We propose a Recruitment of active OSS maintainers to participate in a vignette study[65]. This would present anonymized contributor profiles with varying metric scores, and measure time‐to‐decision (vet vs. reject) and how variations on the proposed ARMS signals would influence these choices. This method would gather qualitative feedback on signal clarity and usefulness.
  2. 2.
    Chilling Effect: A potential issue with establishing an actor reputation system is the possibility of “chilling effect,”[66] where contributors may hesitate or reduce participation if they know their interactions are being tracked and scored [66, 67, 68]. We propose a user survey study to assess contributors’ willingness to participate under different tracking scenarios (e.g., “all projects” vs. “critical only”), and to quantify perceived privacy risks and impact of monitoring activities on contribution intent.
    Ultimately, we do expect some chilling effects. To mitigate them, we recommend limiting the deployment of an ARMS approach to high-impact, security-sensitive OSS projects (e.g., the Linux kernel) rather than applying it to hobby or low-risk repositories. To effectively do this, there is need to develop an effective project importance score. OpenSSF’s criticality score [69] may be used or serve as a starting point. We recommend surveying OSS contributors to gauge a suitable threshold of criticality.
V-C Worked Examples
We illustrate how our proposed system could have worked in the earlier described examples section III-B.
V-C1 XZ Utils Attack
@@
The timeline of events leading to the XZ Utils backdoor reveals several characteristics that map to our proposed security signals [70]:
  1. 1.
@@
  3. 3.
    Targeted feature change requests: Their first public change request focused on adding features to a small set of projects rather than fixing issues.
Under our framework, these traits would yield a low reputation score:
  • Signal Metrics (Signals 1-7) penalize sparse or opaque contribution histories.
@@ V-C2 Dexcom
V-C3 ESLint Compromise
The 2018 ESLint compromise was traced to an attacker breaching a maintainer’s account—enabled by the absence of two-factor authentication. This type of attack cannot be modeled by reputation systems because the attacker took over a legitimate account without performing any genuine contributor actions or exhibiting the behavioral signals that these systems track. This is why we deem it out of scope in the system threat Model section III.
VI Discussions and Future Works
VI-A Threats to Validity
@@
  1. 1.
    False Positives and Negatives: There is a risk of mis-assigning reputation, resulting in inappropriate characterizations of users as more or less trustworthy. Our definitions and metrics do not account for all interactions that could detect incompetence or malicious activities, especially non-artifact interactions such as social engagements, organizations, security communications, feedback to code reviews, etc. To mitigate this, we recommend a weighted combination of metrics, with each metric’s effectiveness considered in section V-A. Using ecosystem-wide averages and standard deviations can also help raise the threshold for issuing advisories based on user scores.
  2. 2.
@@
    Gameability of Proposed Metrics: Our security metrics, like any trust-based system, can be exploited. Users may act, individually or in collusion, to enhance their reputation. We mitigate this risk with time-weighted scoring and recommend incorporating human oversight (e.g., reporters and moderators) for nuanced judgments. However, we acknowledge that trust and safety issues plague all online platforms [71], and a reputation system would be no exception.
  6. 6.
    Possibility of Chilling Effect: Discussed in section V-B.
VI-B Evaluating Actor Intent
@@
Our work assumes that actors are well-intentioned and attributes security failures to genuine contributors’ lack of expertise or mistakes; it does not account for malicious intent. However, some supply-chain incidents arise from malicious actors using techniques such as account spoofing or credential theft. Incorporating intent into reputation systems substantially increases their complexity. Although our current security signals focus on expertise and behavior, they do not capture actor intent. Future work should extend these metrics to infer intent— for example, by analyzing anomalous contribution patterns, unusual social graph connections, or timing irregularities, ultimately creating a more comprehensive reputation model.
VI-C Supporting Ecosystem Heterogeneity
@@
Reputation systems fundamentally face actor‐identification challenges [72, 73]. This issue is especially pronounced in open-source ecosystems, where contributors manage artifacts across multiple platforms—from source repositories to package registries. Although next-generation software-signing tools have improved artifact-to-actor verification [4, 74, 75], reliable cross-platform identification remains vulnerable to identity theft, impersonation, and other attacks. Consequently, ARMS requires stronger mechanisms to verify and vouch for identities across diverse environments.
@@
To provide possible solutions, recent initiatives (e.g., CISA’s RFI for software identifier ecosystems [76]) propose establishing identities through multiple institutions—some designed to preserve privacy [77, 78]. Adapting these models across ecosystems introduces the dual challenges of federation—enabling actors in one ecosystem to be recognized in another—and roaming—allowing actors to transfer their identifier and reputation between providers. Federation and Roaming are challenges reminiscent to other federation protocols work (e.g., OIDC [79] and Distributed Identities [80]). Thus, for an actor reputation system like ARMS, institutions across ecosystems must collaborate and share reputation information whenever a software artifact from one ecosystem A is used in another ecosystem B.
VI-D Social & Process Signals: Beyond Artifact Contributions
@@
Our current security (and weightage) signals (Table I) focus exclusively on an actor’s artifact contributions. As noted in our approach criticisms, we have not yet captured an actor’s interactions with other users—such as code-review feedback, issue discussions, and peer endorsements—which could provide valuable reputational insights.
@@
We also recommend developing process-based metrics to evaluate informal OSS practices. Process-based methods [81] are most effective in structured development environments. However, in open-source ecosystems—where formal processes are often absent [82]—proxy process metrics can still gauge adherence to security best practices. For example, one could measure the frequency of explicit threat-model or design-review discussions in issues or change requests, track the presence and pass rate of automated security scans (e.g., SAST, SBOM generation) in continuous integration workflows, and monitor the regularity of dependency updates following published vulnerability disclosures. By combining artifact, social, and process signals, ARMS can deliver a more holistic reputation assessment for contributors in open-source ecosystems.
VI-E Ethical and Privacy Considerations

### v2

@@ 4.2. Interaction Data – Security Signals and Metrics Definitions
Our system computes reputation from an actor’s historical interactions in the ecosystem by operationalizing them as seven security signals (S1–S7), each derived from one or more measurable metrics (Table 1). We compute each signal score by aggregating its associated metrics (e.g., S1 aggregates three metrics).
@@
    Security standards and recommendations: We consulted frameworks like the SLSA security framework (The Linux Foundation, 2023), the CNCF software supply chain security guidelines (Cloud Native Computing Foundation, 2021), NIST SSDF (Souppaya et al., 2022), NIST SP 800-204D (Chandramouli et al., 2024), OpenSSF S2C2F (OpenSSF, nd), and the CIS Software Supply Chain Security Guide (Center for Internet Security, 2021). We use common recommendations across these sources to prioritize well-established practices.
  2. (2)
    Available security tools in the OSS ecosystem: We identified security tools available through GitHub’s user interface and API, which reflect the security capabilities easily available to contributors on the platform. ARMS does not propose new vulnerability detection or scanning techniques; it consumes best-available ecosystem outputs (e.g., alerts, advisories, and scanner results) and treats them as noisy evidence for decision support rather than ground truth.
Scope and attribution. To avoid conflating actor behavior with project governance, we distinguish actor-behavior signals from repository-governance context and ecosystem context, as described in Table 1. In addition to the seven security signals, our model includes three weighting (impact) signals (W1–W3) that calibrate risk exposure and uncertainty (package usage, community tenure, and centrality). These weighting signals are applied in the reputation computation (section 4.3) to calibrate exposure and uncertainty, rather than being treated as additional security signals.
4.3. Trust Engine – Reputation Computation
In this section, we describe candidate approaches for computing reputation from the signals in Table 1 (see Figure 1). We focus on how interaction histories can be translated into per-signal evidence and a composite score that supports maintainer triage. Because multiple design choices are plausible, we present one concrete instantiation for clarity and identify alternatives that we plan to compare empirically (section 5.1).
First, the interaction data formatter extracts each contributor’s ecosystem interaction history and computes metric values for all proposed metrics in Table 1. Next, the security signal scorer aggregates those metric values into seven security signal scores (S1–S7), where each signal score is computed from its associated metrics (e.g., Signal 1 aggregates three proposed metrics; other signals aggregate one or more metrics).
Signal scoring. For each security signal $s\in\{S1,\dots,S7\}$, one candidate approach is to compute a normalized score $\mathrm{Score}_{s}(a)\in[0,1]$ for actor $a$ over a time window, optionally using time decay to down-weight older interactions. When a metric reflects repository configuration rather than an actor’s direct actions, we treat it as governance context unless control can be attributed to the actor (e.g., the actor is an owner/maintainer or the configuration change is attributable to the actor). Other alternative scoring functions are plausible (e.g., robust aggregation such as medians, or cohort-normalized scoring).
Composite scoring. A simple baseline is to aggregate metric evidence across contribution events $e\in\mathcal{E}(a)$ into each signal score, calibrating each event by package-usage exposure (W1):
$\mathrm{Score}_{S_{s}}(a)=\mathrm{Norm}\!\left(\sum_{e\in\mathcal{E}(a)}\mathrm{W1}(e)\cdot f_{s}(e)\right),$
where $f_{s}(e)$ extracts the event-level contribution to signal $s$ (computed from the proposed metrics in Table 1), and $\mathrm{Norm}(\cdot)$ maps scores to $[0,1]$. Other aggregation strategies may be preferable in practice, including time-decayed sums, winsorized/trimmed statistics, or per-metric normalization before pooling; we treat these as design choices to be evaluated.
Given per-signal scores, one candidate composite is a weighted linear combination:
$R_{0}(a)=\sum_{s=1}^{7}\alpha_{s}\cdot\mathrm{Score}_{S_{s}}(a).$
Combining multiple signals can reduce reliance on any single noisy indicator and mitigate false positives and false negatives, similar in spirit to composite security assessments such as OpenSSF Scorecard (Zahan et al., 2023). We treat the selection of $\alpha_{s}$ as an empirical question: one approach is to set $\alpha_{s}$ via effectiveness analysis (section 5.1), but alternatives include uniform weights, expert-set weights, or learned weights from historical outcomes.
Terminology. We use $\alpha_{s}$ to denote signal importance weights that determine how strongly each of the seven security signals contributes to the composite score. We reserve W1–W3 for calibration factors: W1 is an exposure factor applied to event-level evidence, while W2 and W3 calibrate the composite score for uncertainty (tenure) and ecosystem context (centrality), as described next.
Impact Score (calibration factors W1–W3 in Table 1): We plan to apply calibration factors at two stages.
(1) Per-contribution exposure calibration (W1): We incorporate package usage (W1) when aggregating event-level evidence into each security signal score. Interactions tied to projects with minimal downstream usage contribute less to signal evidence because they present lower supply chain exposure. Concretely, we calibrate contribution events by a bounded usage factor and then aggregate to compute each $\mathrm{Score}_{S_{s}}(a)$.
(2) Composite-level calibration (W2–W3): After computing the seven security signal scores, another approach is to calibrate the composite reputation using community tenure (W2) and centrality (W3). Tenure can moderate overconfidence from sparse histories, while centrality can provide ecosystem context based on connectedness and the time taken to establish it. One candidate formulation is:
$R_{1}(a)=R_{0}(a)\cdot\mathrm{Calibrate}(\mathrm{W2}(a),\mathrm{W3}(a)),$
where $R_{0}(a)=\sum_{s=1}^{7}\alpha_{s}\cdot\mathrm{Score}_{S_{s}}(a)$. Alternative calibration strategies include separating a score from an explicit confidence estimate, using additive rather than multiplicative adjustments, or applying cohort-based calibration.
Benchmark Score: Finally, we consider contextualizing $R_{1}(a)$ relative to ecosystem-wide behavior to support interpretable decisions. One approach is to compute a benchmarked reputation using either an ecosystem percentile or a standardized score relative to the ecosystem median/mean:
$R(a)=\mathrm{Benchmark}(R_{1}(a)).$
We treat benchmarking choice as another design decision within the midst of other alternatives; percentile-, standardized-, and cohort-based baselines (section 5.1).
4.4. Actionability Output: Using ARMS in Maintainer Workflows
ARMS is a decision-support tool that guides triage and review efforts rather than automatically accepting or rejecting changes. For sparse histories, ARMS recommends additional verification (e.g., maintainer review, smaller scoped PRs, CI pass) rather than rejection. For low scores in high-exposure contexts, ARMS recommends stronger safeguards (e.g., two-person review, restrict to non-critical modules, require signing where feasible). For high scores, ARMS recommends faster routing while preserving normal repository policies. ARMS should also surface brief explanations (top contributing signals) and treat missing signals as unknown rather than negative evidence.
4.5. Worked Examples
We illustrate ARMS with the motivating examples from section 3.2.
4.5.1. XZ Utils Attack
@@
The timeline of events leading to the XZ Utils backdoor reveals several characteristics that map to our proposed security signals (Smith, 2024):
  1. (1)
@@
  3. (3)
    Targeted feature pull requests: Their first public pull request focused on adding features to a small set of projects rather than fixing issues.
Under our framework, these traits would yield a low reputation:
  • Signals S1–S7 yield limited positive evidence due to sparse or opaque public contribution history.
@@ 4.5.2. Dexcom
4.5.3. ESLint Compromise
The 2018 ESLint compromise was traced to an attacker breaching a maintainer’s account—enabled by the absence of two-factor authentication. This type of attack cannot be modeled by reputation systems because the attacker took over a legitimate account without performing any genuine contributor actions or exhibiting the behavioral signals that these systems track. This is why we deem it out of scope for our threat Model section 3.
5. Design of Experiments for ARMS
For a first journal submission (TSE), we will prioritize an ARMS prototype and quantitative feasibility plus retrospective sanity checks of the proposed signals (scalability, behavior on well-established maintainers, and known-failure catalogs). We defer user behavioral studies (vignettes, surveys, chilling-effect analysis) to follow-up work after quantitative validation.
5.1. Security Metrics Effectiveness Evaluation – Quasi-Experimental Analyses
To evaluate the utility of the proposed security signals, we propose quasi-experimental studies that test whether these signals (and alternative scoring choices from §4.3) predict downstream security outcomes and observable security-relevant behaviors.
  1. (1)
    Security Practice Adoption Study (Difference-in-Differences): Compare OSS projects that adopted a given security practice (e.g., branch protection, private vulnerability reporting) to matched control projects that did not, using a difference-in-differences design. We will analyze the results with regression models incorporating project and time fixed effects, and (when appropriate) interaction terms for paired practices to isolate combined impacts on security outcomes.
    • Data: Historical GitHub records for projects within the same domain and with similar size and activity levels.
    • Outcomes: Changes in vulnerability incidence and remediation behavior (Signals S1–S2) and shifts in security-relevant workflows (Signals S3–S7).
    • Ethics: Uses only publicly available data.
  2. (2)
    Retrospective Incident Prediction Study: Identify projects that experienced known supply-chain incidents (e.g., npm/Event-Stream, ua-parser-js) and compare their pre-incident signal profiles against matched projects without known incidents.
    • Data: Metric values for each project and its maintainers in a fixed pre-incident window.
    • Analysis: Logistic regression (including interaction terms where justified) to estimate which signals and combinations are most predictive of incident occurrence.
    • Outputs: Predictive utility summaries such as AUC, calibration curves, and error rates under different thresholds.
  3. (3)
    Ablation and Sensitivity Analyses over the Trust Engine Design: Because multiple reputation-computation choices are plausible (§4.3), we will ablate key design decisions and measure how conclusions change under alternative formulations.
    • Signal aggregation: sum vs. median/robust aggregation; with and without time decay.
    • Composite function: linear combination vs. non-linear pooling and thresholded rules that require minimum evidence.
    • Weight selection: uniform weights vs. expert-set weights vs. weights derived from effectiveness analysis.
    • Calibration and benchmarking: multiplicative vs. additive calibration; percentile vs. standardized vs. cohort-based benchmarks.
    • Evaluation: compare predictive performance, stability of rankings, and sensitivity for newcomers (sparse histories) across variants.
5.2. User Behavioural Studies – Surveys & Interviews
  1. (1)
    Collaborator Vetting Study: We propose a Recruitment of active OSS maintainers to participate in a vignette study(Aguinis and Bradley, 2014). This would present anonymized contributor profiles with varying metric scores, and measure time‐to‐decision (vet vs. reject) and how variations on the proposed ARMS signals would influence these choices. This method would gather qualitative feedback on signal clarity and usefulness.
  2. (2)
    Chilling Effect: A potential issue with establishing an actor reputation system is the possibility of a “chilling effect,”(Marder et al., 2016) where contributors may hesitate or reduce participation if they know their interactions are being tracked and scored (Marder et al., 2016; Büchi et al., 2022; Bernauer and Dapp, 2009). We propose a user survey study to assess contributors’ willingness to participate under different tracking scenarios (e.g., “all projects” vs. “critical only”), and to quantify perceived privacy risks and impact of monitoring activities on contribution intent.
    Ultimately, we do expect some chilling effects. To mitigate them, we recommend limiting the deployment of an ARMS approach to high-impact, security-sensitive OSS projects (e.g., the Linux kernel) rather than applying it to hobby or low-risk repositories. To do this, there is a need to develop an effective project importance score. OpenSSF’s criticality score (Arya et al., 2023) may be used or serve as a starting point. We recommend surveying OSS contributors to gauge a suitable threshold of criticality.
6. Discussions and Future Works
6.1. Threats to Validity
@@
  1. (1)
    False Positives and Negatives: There is a risk of mis-assigning reputation, resulting in inappropriate characterizations of users as more or less trustworthy. Our definitions and metrics do not account for all interactions that could detect incompetence or malicious activities, especially non-artifact interactions such as social engagements, organizations, security communications, feedback to code reviews, etc. To mitigate this, we recommend a weighted combination of metrics, with each metric’s effectiveness considered in section 5.1. Using ecosystem-wide averages and standard deviations can also help raise the threshold for issuing advisories based on user scores.
  2. (2)
@@
    Gameability of Proposed Metrics: Our security metrics, like any trust-based system, can be exploited. Users may act, individually or in collusion, to enhance their reputation. We mitigate this risk with time-weighted scoring and recommend incorporating human oversight (e.g., reporters and moderators) for nuanced judgments. However, we acknowledge that trust and safety issues plague all online platforms (Cramer et al., 2025), and a reputation system would be no exception.
  6. (6)
    Possibility of Chilling Effect: Discussed in section 5.2.
6.2. Evaluating Actor Intent
@@
Our work assumes that actors are well-intentioned and attributes security failures to genuine contributors’ lack of expertise or mistakes; it does not account for malicious intent. However, some supply-chain incidents arise from malicious actors using techniques such as account spoofing or credential theft. Incorporating intent into reputation systems substantially increases their complexity. Although our current security signals focus on expertise and behavior, they do not capture actor intent. Future work should extend these metrics to infer intent, e.g., by analyzing anomalous contribution patterns, unusual social graph connections, or timing irregularities, ultimately creating a more comprehensive reputation model.
6.3. Supporting Ecosystem Heterogeneity
@@
Reputation systems fundamentally face actor‐identification challenges (Swamynathan et al., 2010; Marti and Garcia-Molina, 2003). This issue is especially pronounced in open-source ecosystems, where contributors manage artifacts across multiple platforms, from source repositories to package registries. Although next-generation software signing tools have improved artifact-to-actor verification (Newman et al., 2022; Kalu et al., 2025; Schorlemmer et al., 2025), reliable cross-platform identification remains vulnerable to identity theft, impersonation, and other attacks. Consequently, ARMS requires stronger mechanisms to verify and vouch for identities across diverse environments.
@@
To provide possible solutions, recent initiatives (e.g., CISA’s RFI for software identifier ecosystems (Center for Internet Security, 2021)) propose establishing identities through multiple institutions, some designed to preserve privacy (Microsoft, nd; Vidos.id, nd). Adapting these models across ecosystems introduces the dual challenges of federation, enabling actors in one ecosystem to be recognized in another, and roaming, allowing actors to transfer their identifier and reputation between providers. Federation and Roaming are challenges reminiscent to other federation protocols work (e.g., OIDC (OpenID Foundation, [n. d.]) and Distributed Identities (Bouma, 2024)). Thus, for an actor reputation system like ARMS, institutions across ecosystems must collaborate and share reputation information whenever a software artifact from one ecosystem is used in another ecosystem.
6.4. Extending ARMS to Agent Reputation
As software agents become more popular and are increasingly entrusted with operating on artifacts that move between humans and automated tooling (e.g., code, configurations, packages, and credentials)(Wang et al., 2025), it is natural to consider extending ARMS from human contributor reputation to agent reputation. This extension may be urgent as emerging agent-sharing and agent-registry concepts (Pautsch et al., 2025; Xie et al., [n. d.]; Jia et al., 2025) make it easier to distribute, reuse, and compose third-party agents, thereby increasing the chance that insecure or malicious agents propagate across workflows. In such ecosystems, ARMS could support evidence-based selection by scoring agents and their publishers using registry-visible signals such as signed manifests, provenance and evidence pointers, and lifecycle metadata (e.g., active, deprecated, revoked), rather than relying on popularity alone (Pautsch et al., 2025). At the same time, agent reputation differs from the human scenario we present because agents can evolve rapidly, execute autonomously, and be invoked as dependencies by other agents. Consequently, reputation must be strongly versioned, incorporate lifecycle and revocation signals, and account for feedback loops in agent-to-agent selection, building on recent proposals for agent sharing infrastructure (Pautsch et al., 2025).
6.5. Social & Process Signals: Beyond Artifact Contributions
@@
Our current security signals (S1–S7) and calibration factors (W1–W3) (Table 1) focus exclusively on an actor’s artifact contributions. As noted in our approach criticisms, we have not yet captured an actor’s interactions with other users, such as code-review feedback, issue discussions, and peer endorsements, which could provide valuable reputational insights.
@@
We also recommend developing process-based metrics to evaluate informal OSS practices. Process-based methods (Clancy et al., 2021) are most effective in structured development environments. However, in open-source ecosystems, where formal processes are often absent (Spinellis and Szyperski, 2004), proxy process metrics can still gauge adherence to security best practices. For example, one could measure the frequency of explicit threat-model or design-review discussions in issues or pull requests, track the presence and pass rate of automated security scans (e.g., SAST, SBOM generation) in continuous integration workflows, and monitor the regularity of dependency updates following published vulnerability disclosures. By combining artifact, social, and process signals, ARMS can deliver a more holistic reputation assessment for contributors in open-source ecosystems.
6.6. Ethical and Privacy Considerations